# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ 1.3 Comment on the proofs
Each component of $\widetilde{K}_{H}$ with this metric is an injective metric space.
We conjecture that each component of $\widetilde{K}_{H}$ is CAT$(0)$ if we put the $\ell^{2}$-metric on $K_{H}$, however, this seems much harder to show. This proposition is good enough for the purpose of gaining control on minimal filling of cycles of any length in the 2-skeleton of $\widetilde{K}_{H}$; namely, by a simple procedure (see [Hae21], also [Che00, CCHO14]), one can produce a Helly graph out of $\widetilde{K}_{H}$ which is closely related to the 1-skeleton of $\widetilde{K}_{H}$, and control of filling cycles in Helly graphs was obtained in [CCHO14].
Note that the subcomplex $\widetilde{K}_{H}$ depends on the choice of $H$. Different choices of $H$ give a family of subcomplexes that cover $\Delta_{S}$. If we want to understand the minimal filling of a 6-cycle in $\Delta_{S}$ and this 6-cycle happens to be inside one of these subcomplexes, then we are more or less done. However, this is not the case for all 6-cycles, and actually, we do not have a complete understanding of the minimal filling of all 6-cycles in $\Delta_{S}$. However, for certain types of 6-cycles, we do have an extra argument if they are not contained in one of these subcomplexes, and we will come back to this point in a later part of the discussion.
From a topological viewpoint, the complex $\widetilde{K}_{H}$ arises naturally from the deconing of the associated $A_{n}$-type arrangement of hyperplanes. When $\Delta_{S}$ is 2-dimensional, the complex $\widetilde{K}_{H}$ was considered by Falk in [Fal95] for a different purpose, with a different metric.  Though the description of the complex in Falk’s paper is different as well, it gives the same complex (see Section 2.4). Thus, from now on, we will refer to the subcomplex $\widetilde{K}_{H}$ of $\Delta_{S}$ as the Falk subcomplex associated with $H$.
We also record a curious consequence related to the conjecture of Charney and Davis that the spherical Deligne complex is CAT$(1)$ with a natural piecewise spherical metric. While our intention was to circumvent this conjecture, we do manage to prove that in the $A_{n}$ case, suitable subcomplexes of the spherical Deligne complex are CAT$(1)$.
Proposition 1.17 (=Proposition 5.7).
Suppose $\Lambda$ is the type $A_{n}$ diagram with vertex set $S$. Let $S^{\prime}$ be three consecutive vertices in $\Lambda$. Then the relative Artin complex $\Delta_{\Lambda,\Lambda^{\prime}}$ with the induced Moussong metric on $\Delta_{\Lambda}$ is CAT$(1)$.
Now we turn to the case that $A_{S}$ is of type $H_{3}$, which is the most technical case in this article. Let $\pi:\Delta\to\mathsf{C}_{S}$ and $H$ as before. We can still define $\widetilde{K}_{H}$, though now we are no longer able to metrize it as an injective metric space, or a CAT$(0)$ or conformally CAT$(0)$ space. However, we can get around this issue by the following procedure.
Let $\mathcal{A}$ be a central arrangement of hyperplanes $\{H_{\lambda}\}_{\lambda\in\Lambda}$ in $\mathbb{R}^{n}$, and let $M(\mathcal{A})=\pi_{1}(\mathbb{C}^{n}-\cup_{\lambda\in\Lambda}(H_{\lambda}\otimes\mathbb{C}))$ and $G(\mathcal{A})=\pi_{1}M(\mathcal{A})$. There is a spherical Deligne complex $\mathbb{S}\mathbb{D}_{\mathcal{A}}$ (see Section 2.4) with an action $G\curvearrowright\mathbb{S}\mathbb{D}_{\mathcal{A}}$. The quotient is $S_{\mathcal{A}}$, which is defined to be the unit sphere in $\mathbb{R}^{n}$ with a cell structure induced by the intersection of this sphere with  the hyperplanes $\{H_{\lambda}\}_{\lambda\in\Lambda}$. Take a hyperplane of $S_{\mathcal{A}}$, one can still define the associated Falk subcomplex of $\mathbb{S}\mathbb{D}_{\mathcal{A}}$ in a similar way.
From now on, we take $\mathcal{A}=\{H_{\lambda}\}_{\lambda\in\Lambda}$ to be the collection of reflection hyperplanes in $\mathbb{R}^{3}$ for the canonical action of the type $H_{3}$ Coxeter group on $\mathbb{R}^{3}$, then the associated $G$ is the pure Artin group $PA_{S}$ of type $H_{3}$, and $\mathbb{S}\mathbb{D}_{\mathcal{A}}$ is isomorphic to $\Delta_{S}$.  Denote the standard generators of the Artin groups $A_{S}$ of type $H_{3}$ by $\{a,b,c\}$, and assume $m_{ab}=5$ and $m_{bc}=3$. It turns out the hardest type of 6-cycles in $\Delta_{S}$ is a 6-cycle $\omega$ whose vertices have types alternating between $\hat{a}$ and $\hat{c}$. This corresponds to an equation in $A_{S}$ of form
$w_{ab}w_{bc}w^{\prime}_{ab}w^{\prime}_{bc}w^{\prime\prime}_{ab}w^{\prime\prime}_{bc}=1,$ (1.18)
where $w_{ab},w^{\prime}_{ab}$ and $w^{\prime\prime}_{ab}$ are (possibly different) words in the subgroup of $A_{S}$ generated by $a$ and $b$. Similarly, we define $w_{bc},w^{\prime}_{bc}$ and $w^{\prime\prime}_{bc}$. Understanding the minimal filling of $\omega$ is closely related to gaining control over each of the terms in the above equation.
Let $\Lambda^{\prime}$ be a subset of $\Lambda$. This gives a sub-arrangement $\mathcal{A}^{\prime}=\{H_{\lambda}\}_{\lambda\in\Lambda^{\prime}}$. The natural inclusion $M(\mathcal{A})\to M(\mathcal{A}^{\prime})$ gives a surjective homomorphism $PA_{S}\to G(\mathcal{A}^{\prime})$. The rough plan is as follows: Suppose we want to control the term $w_{ab}$ in the above equation. We first choose carefully the subset $\Lambda^{\prime}$ such that the quotient $PA_{S}\to G(\mathcal{A}^{\prime})$ restricts to an injective map on the pure Artin subgroup $PA_{ab}$. This enables us to consider the same equation in the quotient $G(\mathcal{A}^{\prime})$ and the term $w_{ab}$ does not disappear under the quotient. Now we look at an appropriate Falk subcomplex of $\mathbb{S}\mathbb{D}_{\mathcal{A}^{\prime}}$. The hope is that it is now easier to metrize this Falk subcomplex as a non-positive curved metric space (again, one has to choose $\Lambda^{\prime}$ carefully for this to happen). Now we can use the non-positive curvature of the Falk subcomplex to fill in 6-cycles in this subcomplex, which can be translated to an estimate of the term $w_{ab}$.
There are a few other important points to successfully implement this plan. The first is that when we make the quotient, we not only want the term $w_{ab}$ to be more or less unchanged, we also want to make sure that the other terms are still “non-trivial” in an appropriate sense - for example, if most of the other terms disappear under the quotient, then our plan would fail. This places extra constraint on $\Lambda^{\prime}$, as a consequence, it makes the task of arranging the non-positive curvature of the Falk subcomplex harder. However, it turns out that we can arrange $\Lambda^{\prime}$ such that the Falk subcomplex has a natural graph of spaces structure, and the term we want to control corresponds to the geometric properties of a vertex space of this graph of spaces. So it suffices to metrize the vertex space so it has non-positive curvature, which makes things easier.
In both cases, $A_{n}$ and $H_{3}$, the above argument only applies to 6-cycles that are contained in a Falk subcomplex, or whose shadows under the above quotienting process are contained in a Falk subcomplex. This can be combined with another ingredient, namely certain types of projections/retractions in $A_{S}$ and $\Delta_{S}$, to study more general 6-cycles. We can think of (1.18) as a loop in the Cayley graph of $A_{S}$, such that each term of the equation corresponds to a subpath of the loop that is contained in a left $A_{ab}$-coset or a left $A_{bc}$-coset. Suppose the first term $A_{ab}$ corresponds to the identity coset $A_{ab}$ in $A_{S}$, and there is a (not necessarily group theoretic) retraction $r:A_{S}\to A_{ab}$ such that the $r$-image of other cosets are “small” in $A_{ab}$, then we will gain control of the term $A_{ab}$.  There are several kinds of retraction $r:A_{S}\to A_{ab}$ defined in the literature, we use the one in [GP12b, CP14, BP23] as it is the easiest to carry out a very explicit computation of the retraction image. However, the retraction images of other cosets onto $A_{ab}$ are not always small enough, so sometimes such retraction does not give useful information. This is more likely to happen if the 6-cycle is contained in a Falk subcomplex. So our method would treat those types of 6-cycles that are either contained in a Falk subcomplex, or behave well under retraction.
A propagation ingredient
In Section 8, we use the following idea to understand minimal fillings of $n$-cycles in the relative Artin complex $\Delta_{S,S^{\prime}}$, where the ambient Artin group $A_{S}$ is not necessary spherical. The previous discussion relies heavily on the fact that the ambient Artin group is spherical, hence does not apply. When $S^{\prime}\subset S$ is spherical, the complex $\Delta_{S,S^{\prime}}$ is again a union of “apartments” which are round spheres. However, again, there is a way of making $\Delta_{S,S^{\prime}}$ “non-positively curved”. For different choices of $S^{\prime}$ in $S$, the complex $\Delta_{S,S^{\prime}}$ is always quasi-isometric to $\Delta_{S}$ (as long as it is connected). Though for one choice of $S^{\prime}$, $\Delta_{S,S^{\prime}}$ could be a union of round spheres; for a different choice $\Delta_{S,S^{\prime}}$ could be a union of Euclidean spaces or hyperbolic spaces.  Actually, as long as $S$ is not spherical, there is always a choice of $S^{\prime}\subset S$ such that the latter possibility happens. This gives some hope to use the geometry of one relative Artin complex to study the geometry of another relative Artin complex, though the naive quasi-isometries between them are useless for understanding the fine local combinatorics of minimal filling disks for short cycles.
Given an $n$-cycle in $\Delta_{S,S^{\prime}}$ with $S^{\prime}$ spherical, under certain extra assumptions, we are able to enlarge $S^{\prime}$ to $S^{\prime\prime}$ such that $\Delta_{S,S^{\prime\prime}}$ can be endowed with an injective metric or a metric with convex geodesic bicombing. Hence, cycles in $\Delta_{S,S^{\prime}}$ are contained in a non-positive curved space $\Delta_{S,S^{\prime\prime}}$ and we can use the appropriate combinatorial counterpart of the non-positive curvature geometry to understand the fine combinatorics of the minimal filling disks of these cycles. Showing $\Delta_{S,S^{\prime\prime}}$ is non-positive curved relies on checking that certain short cycles in the vertex links have minimal fillings in the links with very specific combinatorial types. However, links of vertices are again relative Artin complexes, with the ambient Artin groups becoming smaller, so this reduces the complexity of the problem.
Ending remarks
The methods discussed here are based on non-positive curvature geometry. In the companion article [Hua24], we use different methods to treat certain cycles in the type $D_{n}$ case. In this whole discussion, we have tried to avoid the more technical point of exactly what kind of 6-cycles we are considering. We end this discussion by mentioning two conjectures, namely Conjecture B.1 and Conjecture B.2, which are about precisely what types of 6-cycles we hope to understand in Artin complexes of spherical type, and what kind of minimal fillings we would expect. After the conjectures, there is also a summary of exactly what kind of 6-cycles we have treated so far. The interested reader is invited to Section A and Section B for a list of conjectures that could potentially extend the results obtained here, as well as some difficulties of the strategy in [Hua23] that remain to be resolved.
1.4 Structure of the article and reading guide
The article has three parts. Part 1 consists of  Section 2 and Section 3, and it is a collection of preliminaries. Part 2 consists of Section 4 to Section 7, where we discuss minimal filling of 6-cycles in different types of spherical Artin complexes. More precisely, Section 4 is about the type $F_{4}$ Artin complex. Section 5 is about a specific type of 6-cycles in type $A_{n}$ and $D_{n}$ Artin complexes. Section 6 and Section 7 are about the type $H_{3}$ Artin complex. Part 3 consists of the remaining sections, where we deduce new cases of the $K(\pi,1)$-conjecture.
All the sections in Part 2 of the article are independent from each other, apart from that  Section 7 relies on Section 6. So after reading Part 1, the reader can take any section in Part 2 and start reading without interruption, with the only exception that reading Section 7 requires material from Section 6.
The main results of Part 2 are Proposition 4.2, Theorem 5.6, and Theorem 7.1. Alternatively, after reading Part 1, if the reader is willing to take these three main results of Part 2 for granted, then the reader can go to Part 3 directly, to see how these three results imply new cases of the $K(\pi,1)$-conjecture. Here is a more detailed description of Part 3. In Section 8 we prove some partial results on how certain properties of 6-cycles in spherical Artin complexes propagate to more general Artin complexes. Section 8 is also independent from Part 2. In Section 9 to Section 11, we use the results from Part 2 and Section 8 to prove all the main theorems.
1.5 Acknowledgment
The author thanks Mike Davis for many long and helpful discussions on hyperbolic type Coxeter groups and Artin groups, as well as the topology of associated hyperplane arrangement complements. The author also thanks Mike Davis for helping with the proof of Corollary 9.3. The author thanks Ruth Charney and Piotr Prytycki for reading part of the article carefully, and giving valuable comments. The author thanks Thomas Haettel for several inspiring discussions on injective metric spaces and Artin groups. The author thanks Damian Osajda and Piotr Prytycki for several inspiring discussions on relative Artin complexes and forms of non-positive. The author also thanks Grigori Avramidi, Mladen Bestvina, Srivatsav Kunnawalkam Elayavalli, Michael Falk, Alexandre Martin, Jon McCommand, Katherine Goldman, Nima Hoda, Camille Horbez, Nick Salter, Rachel Skipper, Antoine Song, and Frank Wagner for helpful related conversation and support.
The author is partially supported by a Sloan fellowship and NSF DMS-2305411. The author thanks the Centre de Recherches Mathématiques in Montreal for the thematic program Geometric Group Theory in 2023 where part of the work was done. The author thanks the organizers of this program. The author thanks the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université) for the trimester program Groups acting on fractals, Hyperbolicity and Self-similarity in 2022 (through LabEx CARMIN, ANR-10-LABX-59-01) where part of the work was done. The author thanks the organizers of this program.
2 Preliminary I: Hyperplane arrangements and related complexes
2.1 Artin groups
Let $\Gamma$ be a finite simplicial graph with each edge labeled by an integer $\geq 2$, with its vertex set $S$. The Artin group (or Artin-Tits group) with presentation graph $\Gamma$, denoted $A_{\Gamma}$ or $A_{S}$, is a group whose generating set is $S$, and there is a relation of the form $aba\cdots=bab\cdots$ with both sides being alternating words of length $m$ whenever two vertices $a$ and $b$ are connected by an edge labeled by $m_{ab}$. Given an Artin group with presentation graph $\Gamma$, its Dynkin diagram $\Lambda$ is obtained from $\Gamma$ by removing all open edges of $\Gamma$ labeled by $2$, and add extra edges labeled by $\infty$ between vertices of $\Gamma$ which are not adjacent. We will also write $A_{\Gamma}$ as $A_{\Lambda}$. The Coxeter group with presentation graph $\Gamma$ (or with Dynkin diagram $\Lambda$), denoted $W_{\Gamma}$ or $W_{\Lambda}$ or $W_{S}$, has the same generating sets and the same relators as the Artin group, with extra relations $v^{2}=1$ for each vertex $v\in\Gamma$. There is a homomorphism $A_{\Gamma}\to W_{\Gamma}$, whose kernel is called the pure Artin group, and is denoted by $PA_{\Gamma}$.
There is a set theoretic section $s:W_{\Gamma}\to A_{\Gamma}$ to the quotient map $A_{\Gamma}\to W_{\Gamma}$ defined as follows. Take an element $g\in W_{\Gamma}$ and we write $g$ as a word $w$ in the free monoid $S^{*}$ generated by $S$. We require $w$ is reduced, i.e. it is a minimal length word in $S^{*}$ representing $g$. Then $s(g)$ is defined to be $w$, viewed as an element in $A_{\Gamma}$. This map is well-defined, as two different reduced words in $S^{*}$ representing the same element of $W_{\Gamma}$ differ by a finite sequence of relators in $A_{\Gamma}$ (cf. [Mat64]).
Recall that for any $S^{\prime}\subset S$ generates a subgroup of $A_{\Gamma}$ which is also an Artin group, whose presentation graph is the full subgraph $\Gamma_{S^{\prime}}$ of $\Gamma$ spanned by $S^{\prime}$ [vdL83]. This subgroup is called a standard parabolic subgroup of type $S^{\prime}$. A parabolic subgroup of $A_{\Gamma}$ of type $S^{\prime}$ is a conjugate of a standard parabolic subgroup of type $S^{\prime}$. A parabolic subgroup of $A_{S}$ is reducible if its type $S^{\prime}$ admits a disjoint non-trivial decomposition $S^{\prime}_{1}\sqcup S^{\prime}_{2}$ such that each element in $S^{\prime}_{1}$ commutes with every element in $S^{\prime}_{2}$. If such decomposition does not exist, then the parabolic subgroup is irreducible.
@@ 2.2 Real arrangements and their dual polyhedrons
Recall that a hyperplane arrangement in a real vector space $V=\mathbb{R}^{n}$ is a locally finite collection $\mathcal{A}$ of affine hyperplanes in $V$. Let $\mathcal{Q}(\mathcal{A})$ be the collection of affine subspaces coming from intersections of elements in $\mathcal{A}$ (note that $V$ is considered an element of $\mathcal{Q}(\mathcal{A})$ as it comes from the empty intersection). Note that each point $x$ in $V$ is contained a unique element in $\mathcal{Q}(\mathcal{A})$ which is minimal with respect to containment. This element is called the support of $x$. A fan of $\mathcal{A}$ is a maximal connected subset of $V$ which is made of points with the same support. Each fan is convex and polyhedral. Denote that collection of all fans of $\mathcal{A}$ by $\operatorname{Fan}(\mathcal{A})$. Note that $V$ is a disjoint union of all elements in $\operatorname{Fan}(\mathcal{A})$. For each $U\in\operatorname{Fan}(\mathcal{A})$, choose a point $x_{U}$ in the relative interior of $U$. The partial order on $\operatorname{Fan}(\mathcal{A})$ is defined by $U_{1}<U_{2}$ if $U_{1}$ is contained in the closure of $U_{2}$ and in this case, we also write $x_{U_{1}}<x_{U_{2}}$. Let $b\Sigma_{\mathcal{A}}$ be a simplicial complex whose vertices are $\{x_{U}\}_{U\in\operatorname{Fan}(\mathcal{A})}$, and whose simplices correspond to chains of form $x_{U_{1}}<x_{U_{2}}<\cdots<x_{U_{k}}$. There is a piecewise linear embedding of $b\Sigma_{\mathcal{A}}$ as a subset of $V$. Now we assemble simplicies of $b\Sigma_{\mathcal{A}}$ to form another cell complex, which we denote by $\Sigma_{\mathcal{A}}$. The closed cells of $\Sigma_{\mathcal{A}}$ are in one-to-one correspondence with vertices of $b\Sigma_{\mathcal{A}}$. We identify the face of $\Sigma_{\mathcal{A}}$ associated with vertex $x_{U}\in b\Sigma_{\mathcal{A}}$ with the union of all simplices of $b\Sigma_{\mathcal{A}}$ corresponding to chains whose smallest element is $x_{U}$. In this way each vertex of $b\Sigma_{\mathcal{A}}$ can also be regarded as the barycenter of a face of $\Sigma_{\mathcal{A}}$.
Suppose $B\in\mathcal{Q}(\mathcal{A})$. A face $F$ of $\Sigma_{\mathcal{A}}$ is dual to $B$ if $F\cap B=\{b_{F}\}$, where $b_{F}$ denotes the barycenter of $F$. Two faces $F_{1}$ and $F_{2}$ of $\Sigma_{\mathcal{A}}$ are parallel if they are dual to the same subspace in $\mathcal{Q}(\mathcal{A})$: we write $F_{1}\parallel F_{2}$. Parallel classes of faces of $\Sigma_{\mathcal{A}}$ are in one-to-one correspondence with elements in $\mathcal{Q}(\mathcal{A})$. For example, the edges dual to a hyperplane in $\mathcal{A}$ form a parallel class of edges in $\Sigma_{\mathcal{A}}$.
@@
Now we consider the properties of nearest point sets between two faces. The proof of the following lemma is left to the reader (using Lemma 2.1 and [DS87]).
Lemma 2.4.
Let $E$ and $F$ be faces of $\Sigma_{\mathcal{A}}$. Define
$X=\{x\in\operatorname{Vert}E\mid d(x,\operatorname{Vert}F)=d(\operatorname{Vert}E,\operatorname{Vert}F)\}$
and
$Y=\{y\in\operatorname{Vert}F\mid d(y,\operatorname{Vert}E)=d(\operatorname{Vert}E,\operatorname{Vert}F)\}.$
Then
  1. 1.
    there are faces $E^{\prime}\subset E$ and $F^{\prime}\subset F$ such that $X=\operatorname{Vert}E^{\prime}$ and $Y=\operatorname{Vert}F^{\prime}$;
  2. 2.
    $\operatorname{Proj}_{E}(\operatorname{Vert}F)=X$ and $\operatorname{Proj}_{F}(\operatorname{Vert}E)=Y$;
  3. 3.
    $\operatorname{Proj}_{E}|_{\operatorname{Vert}F^{\prime}}$ and $\operatorname{Proj}_{F}|_{\operatorname{Vert}E^{\prime}}$ gives a bijection and its inverse between $E^{\prime}$ and $F^{\prime}$;
  4. 4.
    if $\mathcal{W}(E^{\prime})$ is the collection of hyperplanes in $\mathcal{A}$ dual to an edge in $E^{\prime}$, then $\mathcal{W}(E^{\prime})=\mathcal{W}(F^{\prime})=\mathcal{W}(E)\cap\mathcal{W}(F)$;
  5. 5.
    if $\mathcal{W}(E)=\mathcal{W}(F)$, then $E=E^{\prime}$ and $F=F^{\prime}$.
In the situation of this lemma we will write $E^{\prime}=\operatorname{Proj}_{E}(F)$.
In the situation of Lemma 2.4 (5), we will say $E$ and $F$ are parallel. In this case, the bijection between $\operatorname{Vert}E$ and $\operatorname{Vert}F$ given by $\operatorname{Proj}_{E}|_{\operatorname{Vert}F}$ and $\operatorname{Proj}_{F}|_{\operatorname{Vert}E}$ are called parallel translation between $E$ and $F$.
Definition 2.5.
Parallel faces $F$ and $F^{\prime}$ of $\Sigma_{\mathcal{A}}$ are adjacent if $F\neq F^{\prime}$ and if they are contained in a face $F_{0}$ with $\dim(F_{0})=\dim(F)+1$.
Definition 2.6.
Let $B\in\mathcal{Q}_{\mathcal{A}}$. Let $F$ and $F^{\prime}$ be two adjacent parallel faces of $\Sigma_{\mathcal{A}}$ that are dual to $B$. An elementary $B$-segment, or an $(F,F^{\prime})$-elementary $B$-segment is a minimal positive path from a vertex $x\in F$ to $x^{\prime}=p(x)\in F^{\prime}$, where $p:F\to F^{\prime}$ is parallel translation.

### v3

@@ 3 One side flagness in type $F_{4}$
Let $S=\{s_{1},s_{2},s_{3},s_{4}\}$ be consecutive nodes in the Coxeter diagram $\Lambda$ of the Artin group of type $F_{4}$. Let $S^{\prime}=\{s_{1},s_{2},s_{3}\}$. We consider the relative Artin complex $\Delta_{S,S^{\prime}}$, whose vertex set is endowed with the order inherited from $s_{1}<s_{2}<s_{3}$. Then the vertex set of $\Delta_{S,S^{\prime}}$ forms a upward flag poset.
Proof.
Let $\Lambda_{1}$ be the Coxeter diagram of type $E_{6}$. Let $\{t_{i}\}_{i=1}^{5}$ be the consecutive nodes of the linear subgraph $\Lambda^{\prime}_{1}$ of length $4$ in $\Lambda_{1}$. Let $t$ be the node in $\Lambda\setminus\Lambda_{1}$. Consider the homomorphism $\phi:A_{\Lambda}\to A_{\Lambda_{1}}$ by sending $s_{i}$ to $t_{i}t_{6-i}$ for $i=1,2$, $s_{3}$ to $t_{3}$ and $s_{4}$ to $t$. By [Cri00], $\phi$ is an injective homomorphism, whose image is the fixed subgroup of the automorphism $\sigma$ of $A_{\Lambda_{1}}$ such that $\sigma(t_{i})=t_{6-i}$ for $1\leq i\leq 5$ and $\sigma(t)=t$. We define $P_{i}=A_{\hat{s}_{i}}$, the subgroup of $A$ generated by all generators except $s_{i}$, and $Q_{i}=A_{\hat{t}_{i}}$.
4 Complexes for hyperplane arrangements
We recall the definition of several cell-complexes associated with a given hyperplane arrangement, and mention some of their properties, as a preparation for later sections.
@@ 4.1 Real arrangements and their dual polyhedrons
Recall that a hyperplane arrangement in a real vector space $\mathbb{R}^{n}$ is a locally finite collection $\mathcal{A}$ of affine hyperplanes in $\mathbb{R}^{n}$. Let $\mathcal{Q}(\mathcal{A})$ be the collection of affine subspaces coming from intersections of elements in $\mathcal{A}$ (here $\mathbb{R}^{n}$ itself is considered an element of $\mathcal{Q}(\mathcal{A})$ as it comes from the empty intersection). Each point $x$ in $\mathbb{R}^{n}$ is contained a unique element in $\mathcal{Q}(\mathcal{A})$ which is minimal with respect to containment. This element is called the support of $x$. A fan of $\mathcal{A}$ is a maximal connected subset of $\mathbb{R}^{n}$ which is made of points with the same support. Each fan is convex and polyhedral. Denote that collection of all fans of $\mathcal{A}$ by $\operatorname{Fan}(\mathcal{A})$. Note that $\mathbb{R}^{n}$ is a disjoint union of all elements in $\operatorname{Fan}(\mathcal{A})$. For each $U\in\operatorname{Fan}(\mathcal{A})$, choose a point $x_{U}$ in the relative interior of $U$. The partial order on $\operatorname{Fan}(\mathcal{A})$ is defined by $U_{1}<U_{2}$ if $U_{1}$ is contained in the closure of $U_{2}$ and in this case, we also write $x_{U_{1}}<x_{U_{2}}$. Let $b\Sigma_{\mathcal{A}}$ be a simplicial complex whose vertices are $\{x_{U}\}_{U\in\operatorname{Fan}(\mathcal{A})}$, and whose simplices correspond to chains of form $x_{U_{1}}<x_{U_{2}}<\cdots<x_{U_{k}}$. There is a piecewise linear embedding of $b\Sigma_{\mathcal{A}}$ as a subset of $\mathbb{R}^{n}$. Now we assemble simplicies of $b\Sigma_{\mathcal{A}}$ to form another cell complex, which we denote by $\Sigma_{\mathcal{A}}$. The closed cells of $\Sigma_{\mathcal{A}}$ are in one-to-one correspondence with vertices of $b\Sigma_{\mathcal{A}}$. We identify the face of $\Sigma_{\mathcal{A}}$ associated with vertex $x_{U}\in b\Sigma_{\mathcal{A}}$ with the union of all simplices of $b\Sigma_{\mathcal{A}}$ corresponding to chains whose smallest element is $x_{U}$. In this way each vertex of $b\Sigma_{\mathcal{A}}$ can also be regarded as the barycenter of a face of $\Sigma_{\mathcal{A}}$. For $B\in\mathcal{Q}(\mathcal{A})$, a face $F$ of $\Sigma_{\mathcal{A}}$ is dual to $B$ if $F\cap B=\{b_{F}\}$, where $b_{F}$ denotes the barycenter of $F$.